In 1820 the population of a midwest city was 20,000. By 1860 it had grown to 22,000. If it continues to grow at the same rate, what will the population be in 1892? Give your answer to the nearest whole number.
Answer by checkley75(3666)

40/2000=32/X NOW CROSS MULTIPLY
40X=32*2000
40X=64000
X=64000/40
X=1600 IS THE PROJECTED INCREASE FROM 1860 TO 1892.
THUS THE TOTAL POPULATION IS EXPECTEDTO BE 22000+1600=23600.